i can't figure out the steps to find the zeros of this equation: y=125x^3-8
Answer by stanbon(75887)

figure out the steps to find the zeros of this equation: y=125x^3-8
----
Solve: 125x^3-8 = 0
---
(5x)^2-2^3 = 0
---
Factor:
(5x-2)(25x^2+10x+4) = 0
----
x = 2/5 and x is the zeroes of 25x^2+10x+4 = 0
----
Solve: 25x^2+10x+4 = 0
---
x = [-10 +- sqrt(100-4*25*4)]/50
----
x = [-10 +- sqrt(-300)]/50
------
x = [-10 +- 10sqrt(3)i]/50
------
x = (-1/5)-(1/5)sqrt(3)i pr x = (-1/5)+(1/5)sqrt(3)i
---
Note: There are 3 solutions
-----------------------------------
Cheers,
Stan H.